# A Complete Characterization of Pythagorean Hodograph Preserving Mappings

### v1

@@ 2.1. PH Curves and PH preserving mappings
The projection map $\Pi(x,y)=(x,0)$ sends any PH curve $r(t)=(x(t),y(t))$ to $\Pi\circ r(t)=(x(t),0)$ with $\|(\Pi\circ r)^{\prime}(t)\|^{2}=(x^{\prime}(t))^{2}$, a square. Hence $\Pi$ maps PH curves to PH curves, but $D\Pi=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ has rank 11 everywhere, so $\Pi$ is not PH-preserving in the sense of Definition 5.
$2.$ Main results
@@ 4.1. The Case $m\leq 2$
throughout this section. Polynomials in the variables $x=(x_{1},\dots,x_{n})$ with coefficients in AA form $A[x]$; similarly $K[x]$ denotes polynomials in $x$ with coefficients in $K$.
@@
We say that $F$ is primitive (with respect to $y$) if $\operatorname{cont}_{y}(F)=1$. Every $F\in A[x]$ decomposes (uniquely up to a unit in AA) as $F=\operatorname{cont}_{y}(F)\cdot\operatorname{pp}_{y}(F)$, where $\operatorname{pp}_{y}(F)\in A[x]$ is primitive (the primitive part).
@@
Write $G=\frac{a}{b}\,G_{0}$ with $a,b\in A$ coprime and $G_{0}\in A[x]$ primitive. Since $F$ and $G_{0}$ are primitive, so is $FG_{0}$ by the famous Gauss Lemma [jacobson1985, p. 152], whence $\operatorname{cont}_{y}(FG_{0})=1$.Now $FG=\frac{a}{b}\,(FG_{0})\in A[x]$ implies that $b$ divides the content of $a(FG_{0})$, which equals $a$. Hence $b$ divides $a$ in AA. By coprimality, $b$ is a unit, so $G\in A[x]$. ∎
@@
Factor $a(y)$ in AA as $a(y)=a_{\sf sf}(y)\,a_{\sf sq}(y)^{2}$ with $a_{\sf sf}$ square-free and $a_{\sf sq}\in A$. Absorb $a_{\sf sq}^{2}$ into the square on the right-hand side of (10) by setting
@@
Precomposing $\Phi$ with a fixed linear isomorphism AA on the source with $A^{\top}A=S$ reduces $S$ to $I_{2}$. In particular, after a fixed linear change of coordinates on the source, $G(u)=\lambda(u)^{2}I_{2}$. This is all we can get from PH preservation on lines only. In order to obtain the full statement of Theorem 6, we look at one non-linear PH curve, the Tschirnhausen cubic of Example 3.
@@ 4.2. The higher-dimensional case
The next observation forces the constant symmetric matrix $S$ to be a scalar multiple of the identity by looking at all 22-planes.
@@
Assume $m\geq 2$. Let $S\in\mathbb{R}^{m\times m}$ be symmetric positive definite and suppose that for every 22-dimensional subspace $\Pi\subset\mathbb{R}^{m}$ the restriction $S|_{\Pi}$ is a scalar multiple of the identity on $\Pi$. Then $S=c\,I_{m}$ for some positive $c\in\mathbb{R}$.
@@
By Lemma 23, $G=\lambda^{2}S$ with $S$ constant and symmetric. Fix any 22-plane $\iota\colon\mathbb{R}^{2}\hookrightarrow\mathbb{R}^{m}$. Since PH curves in $\mathbb{R}^{2}$ are PH in $\mathbb{R}^{m}$, the map $\Phi\circ\iota\colon\mathbb{R}^{2}\to\mathbb{R}^{n}$ is PH preserving. Applying the already proved statement of Theorem 6 for the case $m=2$ to $\Phi\circ\iota$ yields $G|_{\Pi}=\lambda^{2}I_{2}$ on $\Pi=\iota(\mathbb{R}^{2})$. But $G|_{\Pi}=\lambda^{2}\,S|_{\Pi}$, hence $S|_{\Pi}$ is a scalar multiple of $I_{2}$ for every 22-plane $\Pi$. By Proposition 24, $S=c\,I_{m}$ with $c\in\mathbb{R}$ positive. Absorbing $\sqrt{c}$ into $\lambda$ yields the desired statement. ∎

### v2

@@ 2.1. PH Curves and PH preserving mappings
The projection map $\Pi(x,y)=(x,0)$ sends any PH curve $r(t)=(x(t),y(t))$ to $\Pi\circ r(t)=(x(t),0)$ with $\|(\Pi\circ r)^{\prime}(t)\|^{2}=(x^{\prime}(t))^{2}$, a square. Hence $\Pi$ maps PH curves to PH curves, but $D\Pi=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ has rank $1$ everywhere, so $\Pi$ is not PH-preserving in the sense of Definition 5.
2.2. Main results
@@ 4.1. The Case $m\leq 2$
throughout this section. Polynomials in the variables $x=(x_{1},\dots,x_{n})$ with coefficients in $A$ form $A[x]$; similarly $K[x]$ denotes polynomials in $x$ with coefficients in $K$.
@@
We say that $F$ is primitive (with respect to $y$) if $\operatorname{cont}_{y}(F)=1$. Every $F\in A[x]$ decomposes (uniquely up to a unit in $A$) as $F=\operatorname{cont}_{y}(F)\cdot\operatorname{pp}_{y}(F)$, where $\operatorname{pp}_{y}(F)\in A[x]$ is primitive (the primitive part).
@@
Write $G=\frac{a}{b}\,G_{0}$ with $a,b\in A$ coprime and $G_{0}\in A[x]$ primitive. Since $F$ and $G_{0}$ are primitive, so is $FG_{0}$ by the famous Gauss Lemma [jacobson1985, p. 152], whence $\operatorname{cont}_{y}(FG_{0})=1$.Now $FG=\frac{a}{b}\,(FG_{0})\in A[x]$ implies that $b$ divides the content of $a(FG_{0})$, which equals $a$. Hence $b$ divides $a$ in $A$. By coprimality, $b$ is a unit, so $G\in A[x]$. ∎
@@
Factor $a(y)$ in $A$ as $a(y)=a_{\sf sf}(y)\,a_{\sf sq}(y)^{2}$ with $a_{\sf sf}$ square-free and $a_{\sf sq}\in A$. Absorb $a_{\sf sq}^{2}$ into the square on the right-hand side of (10) by setting
@@
Precomposing $\Phi$ with a fixed linear isomorphism $A$ on the source with $A^{\top}A=S$ reduces $S$ to $I_{2}$. In particular, after a fixed linear change of coordinates on the source, $G(u)=\lambda(u)^{2}I_{2}$. This is all we can get from PH preservation on lines only. In order to obtain the full statement of Theorem 6, we look at one non-linear PH curve, the Tschirnhausen cubic of Example 3.
@@ 4.2. The higher-dimensional case
The next observation forces the constant symmetric matrix $S$ to be a scalar multiple of the identity by looking at all $2$-planes.
@@
Assume $m\geq 2$. Let $S\in\mathbb{R}^{m\times m}$ be symmetric positive definite and suppose that for every $2$-dimensional subspace $\Pi\subset\mathbb{R}^{m}$ the restriction $S|_{\Pi}$ is a scalar multiple of the identity on $\Pi$. Then $S=c\,I_{m}$ for some positive $c\in\mathbb{R}$.
@@
By Lemma 23, $G=\lambda^{2}S$ with $S$ constant and symmetric. Fix any $2$-plane $\iota\colon\mathbb{R}^{2}\hookrightarrow\mathbb{R}^{m}$. Since PH curves in $\mathbb{R}^{2}$ are PH in $\mathbb{R}^{m}$, the map $\Phi\circ\iota\colon\mathbb{R}^{2}\to\mathbb{R}^{n}$ is PH preserving. Applying the already proved statement of Theorem 6 for the case $m=2$ to $\Phi\circ\iota$ yields $G|_{\Pi}=\lambda^{2}I_{2}$ on $\Pi=\iota(\mathbb{R}^{2})$. But $G|_{\Pi}=\lambda^{2}\,S|_{\Pi}$, hence $S|_{\Pi}$ is a scalar multiple of $I_{2}$ for every $2$-plane $\Pi$. By Proposition 24, $S=c\,I_{m}$ with $c\in\mathbb{R}$ positive. Absorbing $\sqrt{c}$ into $\lambda$ yields the desired statement. ∎